a triangle has a perimeter of 165 cm. the first side is 65cm less than twice the second side. the third side is 10cm less than the second side. write and solve an equation to find the length of each side of the triangle
Answer by MLipsky(9)


FIRST, we need variables. Once we define variables, it's much easier to turn this word problem into a math problem.

Let a = first side
Let b = second side
Let c = third side

NOW, we can turn the words into equations:

"a triangle has a perimeter of 165 cm."
a + b + c = 165

"the first side is 65cm less than twice the second side."
a = 2b - 65

"the third side is 10cm less than the second side."
c = b - 10

Before we finish, I have to ask: Who writes problems like this??? Pointless problems like these are why kids don't like math! Ugh. Drives me crazy. It's a shame, because solving math problems really does have a certain satisfaction, once you learn how. [Okay. I'm done. Back to the problem.]

If we could get this to have only one variable, we could solve it. Substitution to the rescue!
a + b + c = 165 (rewrote equation from above)
(2b - 65) + b + (b - 10) = 165 (substituted "a" and "c" from equations above)

See how that works? Let's solve it.
2b - 65 + b + b - 10 = 165 (dropped the parentheses, because there was nothing to distribute, not even a minus sign)
4b - 75 = 165 (combined like terms)
4b = 240 (added 75 to both sides)
b = 60 (divided both sides by 4)

We found the second side! We can find the first and third sides using those equations from above:
a = 2b - 65
a = 2*60 - 65
a = 120 - 65
a = 55

c = b - 10
c = 60 - 10
c = 50

All done.


For online tutoring, or worked problems by e-mail, please visit
Math Locker.com